Suppose a horizontal wind blows with a speed of 15.4 m/s outside a large pane of plate glass with dimensions 4.00x1.30m. Assume the density of the air to be 1.30 kg/m^3. The air inside the building is atmospheric pressure.
I have already solved and verified the following but need help with (c):

(a) What is the pressure difference on either side of the window pane? 154.15 Pa
(b) What is the net force on the window pane? 801.6 N
(c) What if a second skyscraper is built nearby, the airspeed can be especially high where wind passes through the narrow separation between buildings. Find the pressure with a wind speed TWICE has high. ??? Need help please!
(d) What is the net force on the windowpane if the wind speed is doubled? 3206.4 N
Answer #1


€ al = £502 X 13x (15-4) == 154.154 Pa F - (AP) A = 154,154 * ( 4X1-3) - 801-6008 N opz +(2012 -{x 1-3x (2x155.4)= 616.616 Pa